# On the Performance of Dual-Antenna Repeater Assisted Bi-Static MIMO ISAC

### v2

@@ VI Numerical Results & Discussions
In this section, we quantify the effects of a repeater on the overall ISAC performance. The transmit power at the transmit BS is $1$ Watt. The path loss for the communication channels is modeled using the 33GPP Urban Microcell model. Each of the BSs is equipped with $8$ antennas. The height of the BS, repeaters, and users are $25$, $10$, and $1.5$ meters, respectively, following the models adopted in [Jianan_TWC]. The carrier frequency and bandwidth are set to $1.9$ GHz and $20$ MHz, respectively. At the receive BS, the noise spectral density is 174-174 dBm/Hz, and the noise figure is $9$ dB [Michail_TWC]. The zone area is $500$ square meters, with the BSs being in $(0,250)$ and $(450,250)$; the sensing zone is $(225,250)$. The repeater is located at $(125,250)$. The user locations are randomly generated. The repeater has the same noise power as the BS. The sensing-related parameters follow the Swerling-I model [Wei_ISAC_Channel_Model] and $\left\{\boldsymbol{a}_{{\mathsf{t}}},\boldsymbol{a}_{{\mathsf{r}}}\right\}$} are generated following [Behdad_Ozlem_TWC]. For clutter, i.i.d. Gaussian taps are assumed. Other relevant parameters are also mentioned alongside the simulation results. The inter-BS channel and the residual interference is modeled according to [AC_CRM_TCOM_2024].
@@
Figure 3: CDF of downlink per-user SE with different choices of precoders. Equal power allocation has been employed by the transmit BS.
@@
Fig. 4 illustrates the effects of a repeater on the target detection. For this, the repeater is placed within $100$ meters of the target hot-spot area, and we consider $K=10$ communication users being served by the same BS, whose locations are generated uniformly at random. The slot duration for the GLRT consists of $50$ channel uses, i.e. $\tau=1,2,\ldots,50$. For simplicity, we assume a Gaussian clutter model. Firstly, we observe that the repeater can substantially improve the target detection performance as it amplifies and forwards the echo from the target and thereby improves the received strength of the echo signal at the BS. Specifically, the repeater-aided system attains a high PoD at a lower value of RCS compared to the traditional bi-static detection scenario. An increase in clutter variance degrades the overall detection performance; however, the repeater-assisted system remains robust to it in contrast with the no repeater case.
In Fig. 4 illustrates the involved trade-offs between the different choices of precoders at the transmit BS. The repeater gain is set to $5$ dB above the noise floor. We can readily observe the degradation of the downlink SE for the target-centric precoder, as such a choice adds additional interference to the downlink users due to the dedicated target signals, for both with and without a repeater being present in the system. Communication-centric precoder nullifies that interference. Secondly, we also observe the benefits in the downlink SE due to the repeater.
@@ VII Remarks and Open Questions
This paper presented the initial results and a mathematical framework that facilitates repeater deployment for bi-static ISAC, and reported the benefits that can be obtained by using a repeater in the target hot-spot area. One interesting future study is to incorporate channel estimation within the developed framework and include the time-varying nature of the channels. Furthermore, the impact of multiple repeaters (i.e. swarm-repeaters) in the system, including the effects of interaction between repeaters along with downlink power control, can be interesting to explore. A detailed comparison with other existing technologies, such as RIS, could also be examined as a future study.
-A Proof of Lemma 1
@@
$\sum\limits_{\tau=1}^{\tau_{{\mathtt{L}}}}{\mathbf{y}}_{\mathtt{BS}}^{H}[\tau]\boldsymbol{\Sigma}_{{\mathsf{s}}}^{-1}[\tau]{\mathbf{y}}_{\mathtt{BS}}+{\mathbf{z}}^{H}{\mathbf{Q}}_{\mathcal{H}_{1}}{\mathbf{z}}-2\Re{\left\{{\mathbf{z}}^{H}\mathbf{t}_{\mathcal{H}_{1}}\right\}},$
@@
$-{\underset{{\left\{\alpha,\dot{{\mathbf{c}}}\right\}}}{{\mathrm{min}}}}f(\alpha,\dot{{\mathbf{c}}};\mathcal{H}_{1})=-\sum\limits_{\tau=1}^{\tau_{{\mathtt{L}}}}{\mathbf{y}}_{\mathtt{BS}}^{H}[\tau]\boldsymbol{\Sigma}_{{\mathsf{s}}}^{-1}[\tau]{\mathbf{y}}_{\mathtt{BS}}+\mathbf{t}_{\mathcal{H}_{1}}^{H}{\mathbf{Q}}_{\mathcal{H}_{1}}^{-1}\mathbf{t}_{\mathcal{H}_{1}}.$ (18)
@@
$={\mathtt{C}}_{\mathcal{H}_{0}}\exp{\left\{\mathbf{t}_{\mathcal{H}_{0}}^{H}{\mathbf{Q}}_{\mathcal{H}_{0}}^{-1}\mathbf{t}_{\mathcal{H}_{0}}-\sum\limits_{\tau=1}^{\tau_{{\mathtt{L}}}}{\mathbf{y}}_{\mathtt{BS}}^{H}[\tau]\boldsymbol{\Sigma}_{{\mathsf{s}}}^{-1}[\tau]{\mathbf{y}}_{\mathtt{BS}}\right\}},$ (19)
where ${\mathtt{C}}_{\mathcal{H}_{0}}$ is a constant, $\mathbf{t}_{\mathcal{H}_{0}}$ and ${\mathbf{Q}}_{\mathcal{H}_{0}}$ are respectively given by (11b) and (11c). Finally, it is easy to show that $\frac{{\mathtt{C}}_{\mathcal{H}_{1}}}{{\mathtt{C}}_{\mathcal{H}_{0}}}=\frac{1}{\pi\sigma^{2}_{{\mathtt{T}}}}$. Finally, combining (18) and (19), we get the final result. ∎

### v3

@@ VI Numerical Results & Discussions
In this section, we quantify the effects of a repeater on the overall ISAC performance. The transmit power at the transmit BS is $1$ Watt. The path loss for the communication channels is modeled using the 33GPP Urban Microcell model. Each of the BSsis equipped with $8$ antennas. The height of the BS, repeaters, and users are $25$, $10$, and $1.5$ meters, respectively, following the models adopted in [1]. The carrier frequency and bandwidth are set to $1.9$ GHz and $20$ MHz, respectively. At the receive BS, the noise spectral density is 174-174 dBm/Hz, and the noise figure is $9$ dB [12]. The zone area is $500$ square meters, with the BSs being in $(0,250)$ and $(450,250)$; the sensing zone is $(225,250)$. The repeater is located at $(125,250)$. The user locations are randomly generated. The repeater has the same noise power as the BS. The sensing-related parameters follow the Swerling-I model [19] and $\left\{\boldsymbol{a}_{{\mathsf{t}}},\boldsymbol{a}_{{\mathsf{r}}}\right\}$} are generated following [17]. For clutter, i.i.d. Gaussian taps are assumed. Other relevant parameters are also mentioned alongside the results. The inter-BS channel and the residual interference are modeled according to [18].
@@
Figure 3: CDF of DL per-user SE with different choices of precoders. Equal power allocation has been employed by the transmit BS.
@@
Fig. 4 illustrates the effects of a repeater on the target detection. For this, the repeater is placed within $100$ meters of the target hot-spot area, and we consider $K=10$ communication users being served by the same BS, whose locations are generated uniformly at random. The slot duration for the GLRT consists of $50$ channel uses, i.e. $\tau=1,2,\ldots,50$. For simplicity, we assume a Gaussian clutter model. Firstly, we observe that the repeater can substantially improve target detection performance as it amplifies and forwards the echo from the target and thereby improves the received echo strength at the BS. Specifically, the repeater-aided system attains a high PoD at a lower value of RCS compared to the traditional bi-static detection scenario. An increase in clutter variance degrades the overall detection performance; however, the repeater-assisted system remains robust to it in contrast with the no repeater case.
In Fig. 4 illustrates the involved trade-offs between the different choices of precoders at the transmit BS. The repeater gain is set to $5$ dB above the noise floor. We can readily observe the degradation of the DL SE for the target-centric precoder, as such a choice adds additional interference to the DL users due to the dedicated target signals, for both with and without a repeater being present in the system. Communication-centric precoder nullifies that interference. We also observe the benefits in the DL SE due to the repeater.
@@
Finally, Fig. 5 manifests the trade-off in the ISAC performance, i.e., DL SE versus PoD/probability of missed detection (PoMD), as the repeater location is varied from a zone where DL users are located towards the target hot-spot area. Specifically, for this, we consider that $4$ downlink users are randomly located within $100$ meters of the transmit BS and then the repeater location (x-axis) is moved from its original location $(125,250)$ towards the target hot-spot area. The RCS value is taken as 30-30 dBsm. The experiment demonstrates that a single repeater can aid communication or target detection based on its location, and thus motivates extending the current framework for a swarm-repeater-assisted system and investigating whether they can potentially offer uniformly good sensing and communication performance.
Figure 5: DL SE versus target-detection (PoD and PoMD) performance trade-off as the repeater location is moved from user locations to target hot-spot area.
@@ VII Remarks and Open Questions
This paper presented initial results and a mathematical framework that facilitates repeater deployment for bi-static ISAC and reported the benefits of using a repeater in the target hot-spot area. One interesting future study is to incorporate channel estimation into the developed framework and to account for the time-varying nature of the channels. Furthermore, the impact of multiple repeaters (i.e. swarm-repeaters) in the system, including the effects of interaction between repeaters along with DL power control, can be interesting to explore.
Finally note that, RISs can be a competing technology with repeaters. However, RIS and repeaters are governed by fundamentally different physics (see [3, Table I] for a detailed comparison). Notably, the repeater is a band-selective device, whereas the RIS reflects all impinging signals regardless of their frequency. Therefore, repeater-assisted systems are not subject to out-of-band interference and are compatible with operational requirements in real-world multi-operator environments. Due to their ease of deployment, repeaters are already being actively considered for standardization in multi-operator systems. The vanilla repeater does not beamform, whereas an RIS can. It remains to be investigated how an $N$-element RIS compares with a swarm of repeaters within a unified mathematical framework.
Appendix A: Proof of Lemma 1
@@
$\sum\nolimits_{\tau=1}^{\tau_{{\mathtt{L}}}}{\mathbf{y}}_{\mathtt{BS}}^{H}[\tau]\boldsymbol{\Sigma}_{{\mathsf{s}}}^{-1}[\tau]{\mathbf{y}}_{\mathtt{BS}}+{\mathbf{z}}^{H}{\mathbf{Q}}_{\mathcal{H}_{1}}{\mathbf{z}}-2\Re{\left\{{\mathbf{z}}^{H}\mathbf{t}_{\mathcal{H}_{1}}\right\}},$
@@
$-{\underset{{\left\{\alpha,\dot{{\mathbf{c}}}\right\}}}{{\mathrm{min}}}}f(\alpha,\dot{{\mathbf{c}}};\mathcal{H}_{1})=-\sum\limits_{\tau=1}^{\tau_{{\mathtt{L}}}}{\mathbf{y}}_{\mathtt{BS}}^{H}[\tau]\boldsymbol{\Sigma}_{{\mathsf{s}}}^{-1}[\tau]{\mathbf{y}}_{\mathtt{BS}}+\mathbf{t}_{\mathcal{H}_{1}}^{H}{\mathbf{Q}}_{\mathcal{H}_{1}}^{-1}\mathbf{t}_{\mathcal{H}_{1}}.$
@@
$={\mathtt{C}}_{\mathcal{H}_{0}}\exp{\left\{\mathbf{t}_{\mathcal{H}_{0}}^{H}{\mathbf{Q}}_{\mathcal{H}_{0}}^{-1}\mathbf{t}_{\mathcal{H}_{0}}-\sum\nolimits_{\tau=1}^{\tau_{{\mathtt{L}}}}{\mathbf{y}}_{\mathtt{BS}}^{H}[\tau]\boldsymbol{\Sigma}_{{\mathsf{s}}}^{-1}[\tau]{\mathbf{y}}_{\mathtt{BS}}\right\}},$
where ${\mathtt{C}}_{\mathcal{H}_{0}}$ is a constant, $\mathbf{t}_{\mathcal{H}_{0}}$ and ${\mathbf{Q}}_{\mathcal{H}_{0}}$ are respectively given by (10b) and (10c). From the above, we get the final result. ∎